Timeline of instrument status
The following table describes the events affecting the quality or the measurement of data obtained with XSHOOTER. Its content can be downloaded as a Comma Separated Values file. It is complete from September 13th, 2020.
The Start Date column indicates when the event took place. Some events cannot be defined by a single instant, but extend over sometimes several days, weeks or months. Operation with the instrument may continue while the event occurs, in which case the Ongoing flag is set. If the event had ended, the Ongoing ? flag is lowered and End Date is given. The Type column indicates the characteristic(s) of the event: Earthquake, Maintenance, Upgrade, Element failure, Intervention, Pipeline installation at QC, Change of nominal values or thresholds, Modification of configuration file of this database, or Other. The Instrument Modes column indicates the affected mode. The column KPI indicates which quality control parameters identified as Key Performance Indicators (i.e., parameters following the most critical quantities) are affected; the column non-KPI indicates which other critical Quality Control parameters are affected. The Calibration Raw Types column indicates the raw types of the affected calibrations. The Calibration Breakpoint flag indicates if calibrations obtained before the event cannot be used for science data obtained after the event or if calibration data obtained after the event cannot be used with science data obtained before the event. The Comment column provides more details. The History column allows one to see possible past information on the entry.
The Quality Control pages for XSHOOTER can be accessed through the QC overview page.
- The IFU is not offerred anymore since ESO Period 111 (April 2023). Users interested in carrying out integral field spectroscopy should consider other alternatives available at the VLT.
- Starting in P108 (October 2021), the convention for acquisitions using blind offsets has changed in order to follow the VLT standard. Please refer to Sect. 3.1.2 of the User manual for more information.
- April 2021: OBs affected by an ADC issue in the period 2017-2019 are now highlighted in the night logs with a comment reporting #ADCISSUE as well as the affected arm. Please download the night log together with your data from the archive to check if there is any such comment in the OB comment field and hence to know if your data was affected. Please note that this is now monitored and no more malfunctioning was reported after December 2019.
- October 2020: X-shooter has been recommissioned at UT3 and science operations have resumed on October 12.
- February 2020: X-shooter is moving from UT2 to UT3 and will be unavailable from 2-31 March 2020. Update: Recommissioning on UT3 is delayed due to the shutdown of Paranal because of the global pandemic.
- January 2020: The ADCs have been occasionally failing in both the UVB and VIS arm since September 2017. Incorrect positions of the ADCs lead to wavelength-dependent slit losses. Affected users have been contacted. We are working on an algorithm to correct the affected data. The malfunction has been corrected on the hardware side and we have implemented software checks to catch potential future failures.
- May 2019: X-shooter will be out of operation between the 7th and 12th of May due to a UT2 Cassegrain intervention.
- August 2018: X-shooter will undergo a maintenance between 29-11-2018 to 7-12-2018 and will be out of operations during this time slot.
- November 2017: In P101, in addition to the change in the calibration plan announced, the Observatory will obtain one telluric for each instrument setting observed during the night. No constraints on airmass and time constraints will be imposed. This applies also to visitor mode observations. The temperature and water vapour profiles from LHATPRO (as measured by the radiometer) allow Molecfit to generate synthetic atmospheric emission and absorption spectra that can be used to efficiently remove telluric features. One limitation is the capability to measure the FWHM of the line spread function, which might be a challenge for low S/N spectra.
- September 2017: Starting with Period 101, the calibration plan does not include observations of telluric standard stars for each observation anymore. Instead, users are encouraged to use Molecfit.
Molecfit can provide accurate telluric line corrections in most cases, provided that a spectral region of about 10 to 40 nm is mostly free of intrinsic spectral features, but is instead affected by telluric absorption lines of low or moderate optical depths caused by water vapor. The signal-to-noise ratio per pixel in this spectral range should be larger than 5 per Observing Block. In the VIS, suggested spectral regions are 810 to 830 nm or 940 to 951 nm. In the NIR, suggested regions are 1300 to 1321 nm, 1745 to 1785 nm, but other regions are possible. However, observation of a telluric standard star is recommended in other cases, such as if
(a) the VIS or NIR spectrum shows low flux in all regions affected by telluric water vapor lines of low or moderate optical depths, so that the signal-to-noise ratio per resolution element is not expected to be larger than 10 per OB in any of these regions
(b) the whole VIS or NIR spectrum is affected by strong intrinsic absorption lines.
The execution time for the observation(s) of the telluric standard star (1 per Observing Block of the science target) must be included in the phase 1 proposal and be submitted in concatenation with the science during phase 2. The Telluric Standard Star Catalogue Search Tool can be used to find appropriate stars (the "Hipparcos/2MASS B stars single" catalogue contains a list of single, B-type, main-sequence stars, verified in previous X-shooter observations).
- September 2017: The Virtual Image Slicer mode allows to observe bright objects that would otherwise saturate, but also to increase the signal to noise ratio in one exposure.
- September 2017: A short report with a characterization and some recommendations regarding X-shooter IFU observations and data analysis can be found here: Characterization of the X-shooter IFU mode.
- May 2017: The new ADC systems are operational. Observations are now obtained by default with working ADCs in the UVB and VIS arms.
- April 2017: X-shooter is unavailable starting April 18th to allow for the installation and commissioning of re-designed ADC systems. We estimate that X-shooter will be back in operation with working ADCs by the end of May.
- Starting in Period 100, we are changing the nighttime calibration plan. The Observatory will supply telluric standard stars only for cases where the requested S/N of the science spectrum in a single OB is 10 or below. For all other cases, if the user requires telluric standard star, observations they must be accounted for in their time request at Phase 1. Users are encouraged to use the molecfit tool, which allows to fit and correct the telluric lines. However, observations of telluric standard stars may be required in some cases (large number of intrinsic features, little or no continuum, low S/N). Users need to consider the following cases: 1.) The requested S/N of the science spectrum in a single OB is >10: no observatory provided telluric standard star. Users that require a telluric need to submit them in a concatenation. The corresponding execution time must be taken into account at Phase 1. 2.) The requested S/N of the science spectrum in a single OB is <10: The user has to explicitly state in the README file that they require a telluric standard star, which then will be provided by the observatory.
- February 2017: VIS readout noise has been recovered. (The increased readout was due to a grounding connection.)
- January 2017: After X-shooter was remounted following the last mirror recoating, the readout noise is much higher in the VIS arm. All observations since 2016-12-20 are affected. Unfortunately, we have not yet identified the cause. The readout noise increased by 20-70% depending on the readout mode used, see http://www.eso.org/observing/dfo/quality/XSHOOTER/reports/HEALTH/trend_report_BIAS_VIS_ron_raw_HC.html.
- December 2016: The use of the MASTER RESPONSE function for the UVB-arm introduces a spectral artefact at λ=365nm, approximately the Balmer discontinuity, in the reduced spectra for data obtained since April 2015. The phase3 science archive contained affected data in 2015-2016.
- August 2015 Spectro-astrometry technique performed with XSHOOTER: Whelan et al. 2015, 579, A48.
- August 2015: A document about the historical wavelength shift between the arms has been released. It presents their origin and the corrective actions that were performed. It is available from the manuals page.
- Two mapping templates (slit and IFU modes) are offered since P96. They facilitate mapping observations and the data reduction with the REFLEX workflow.
- In April 2015, a new pipeline version will be released. It corrects several bugs in the wavelength calibration due to non optimal correction of the flexures leading to offsets of a few km/s. There are still small residuals in the radial velocity depending on the position in the slit, which need to be fully characterized.
- In P96, the ETC will take into account the difference between image quality and seeing.
- Due to a software issue, the target centering was inaccurate in the y direction by 0.25" for g_prime and U band filter acquisitions between Dec 2013 and 2 Sep 2014. The centering was possibly not optimum at some point between 2008 and 2013 for all filters with a shift of 0.17".
- In Period 94, XSHOOTER is offered again at UT2. The instrument will be dismounted from UT3 early September 2014. The re-commissiong of the instrument on UT2 is scheduled for mid-October 2014. X-shooter will be operational starting on October 16.
- Some tests were performed of a diaphragm mode that allows to observe very bright objects by diaphragming the entrance of the instrument. A small report is available here and some reduced data obtained during the technical night are available there.
- Since an upgrade of the VLT SW in January 2014, the quality of the aquisition images is suboptimal (higher bias and noise level). The imaging mode is not affected.
- Oct 2013: XSHOOTER is now installed at UT3. The efficiency of the telescope + instrument is better by 5-10% in all arms compared to UT2.
- In P93, a light imaging mode performed with the fourth channel (A&G camera) will be offered.
- No Large Programs are accepted from P93 onwards to allow for an eventual intervention to repair the ADC drives.
- In P92, XSHOOTER will be moved from the UT2 to the UT3 Cassegrain focus.
- IMPORTANT: ADCs problems. Due to mechanical problems with the ADC (atmospheric dispersion corrector) drives, the ADCs were disabled on August 1st 2012. During the previous months the ADCs for the UVB and VIS arms have been first occasionally and then increasingly failing. Incorrect position of ADCs leads to slit losses worse than if they are not used. An intervention to fix this problem is under evaluation. To minimize the slit losses, we advise to prepare observations as follows:
1) Put the slit at parallactic angle (default value 9999), if possible.
2) Observe the target close to the meridian, at low airmass. The mimimum airmass of your target will depend on its declination. As a reference a plot of the airmass vs. target declination corresponding to observing +/-2h from the meridian is shown on the VIMOS Phase 2 web page: VIMOS visibilities
- A document describing the NIR readout mode, its regimes, and the impacts is available here.
- In July 2011 the NIR slits wheel was replaced by a new one with two new slits of 0.6" and 0.9" with a K-band blocking filter allowing observations not dominated by the thermal background in J and H bands (background reduced by factors 3 to 4). Measurements show that the readout noise is the new limiting factor for these two slits. The 1.5" slit was removed. In addition, the TCCD performance was improved allowing a reduced pick-up noise.
- New templates are available for P88 (spectrophotometric observations in nodding mode, additional AFC during an OB).
- An oscillation of the dichroic absorption band in the UVB and VIS arms was reported. The last UVB arm order and the first four orders of the VIS arm suffer from this oscillation. More information on the Known Problems webpage: (see here)
- October 2010: the pattern problem in the VIS arm has been fixed for both readout modes and the readout noise is constant.
- There is a meeting devoted to the 1st year of XSHOOTER science (October 19-22, 2010): website
- May 2010: XSHOOTER pipeline, current version 1.2.2: here. First public release of XSHOOTER pipeline, v1.0.0.
- Feb 2010: The ESO Archive Group has prepared the XSHOOTER instrument specific query form: here.
- New SM operation requirements/constraints might be implemented each new period. Please, always check the User Manual and the instrument specific phase 2 requirements or phase 2 news.
- In Nov 2008 and Jan 2009 the UVB and VIS arms have been commissioned at the UT3 Cassegrain focus (see Messenger 134, 12 ).
- The NIR arm has been commissioned in March 2009.
- XSHOOTER commissioning data (Nov 2008, Jan 2009, Mar 2009 and May 2009) are public: here.
- Paranalization data (29 May to 1 Jun 2009) are public: here.
- XSHOOTER science verification (SV) data (Aug and Sep 2009) are public: here.